# Actions Speak Louder Than Words: Rate-Reward Trade-off in Markov Decision Processes

### v1

@@ 6 Experimental Results
We evaluate Act2Comm across three distinct MDP environments, as detailed in Appendix D, with communication performance measured by the bit error rate (BER). Due to page limitations, the results of environment, “Erratic robot”, are provided in the Appendix D.4.
Experiment 1: Lucky Wheel. In this game, agent keeps spinning a wheel to accumulate rewards by choosing either clockwise or counterclockwise. It is modeled as an MDP with 3 states and 2 actions, the details of the environment and experimental setting is provided in Fig. 9(a) in Appendix D.
We examine the trade-off among the three performance metrics in Fig. 3. We set the optimal reward-maximizing policy as the target policy, and consider different code rates. By adjusting $\lambda$, we can control how closely the coding policy approximates the target policy. The shaded regions in the figures (a)-(c) represent achievable regions of Act2Comm with various $\lambda$. When $\lambda$ is large, regardless of the coding rate, Act2Comm learns a policy that mirrors the target policy for the optimal reward. In this case, all messages are mapped to the same sequence of decision rules since the target policy is stationary and deterministic. As a result, the BER is 0.5, indicating no communication capability.
@@
Experiment 2: Catch the Ball. Next we consider “Catch the Ball”, which is an MDP with 27 states and 3 actions. This MDP includes a parameter $p$ that influences its transition matrix (see the Appendix D for details). When $p=0$, the action-state channel is perfect as each action can be reliably inferred from the resulting state transition, but it becomes noisy for $p>0$.
@@ Appendix A Notation and Definitions
To bridge the RL and FSC areas, we provide a notation table for this paper. Note that lowercase and uppercase bold letters represent vectors and matrices, respectively.
@@
$\mathcal{X}$, $\mathcal{Y}$, $\mathcal{S}$ Channel input, output alphabet, and state alphabet
@@ B.1 Proof of Theorem 1
    $r(s_{i},q_{i}(s_{i}))=\sum_{x_{i}\in\mathcal{X}}q_{i}(x_{i}|s_{i})\sum_{s_{i+1}\in\mathcal{S}}\bm{T}(s_{i+1}|s_{i},x_{i})\log\frac{\bm{T}(s_{i+1}|s_{i},x_{i})}{\sum_{x^{\prime}_{i}}\bm{T}(s_{i+1}|s_{i},x^{\prime}_{i}){q_{i}}(x^{\prime}_{i}|s_{i})}$
@@
$C=\max_{\{q_{i}\}_{i\geq 1}}\lim_{N\to\infty}\frac{1}{N}\sum_{i=1}^{N}\sum_{s_{i}\in\mathcal{S}}P(s_{i})r(s_{i},q_{i}(s_{i}))$ (16)
@@
    Any stationary policy yields a Markov chain with one ergodic class and a possibly empty set of transient states (Assumption 1).
@@
$C=\max_{\{\pi(\cdot|s)\in\Delta(\mathcal{X}):s\in\mathcal{S}\}}\lim_{N\to\infty}\frac{1}{N}\sum_{i=1}^{N}I(X_{i};S_{i+1}|S_{i},s_{1})$ (17)
Note that each $\{\pi(\cdot|s)\in\Delta(\mathcal{X}):s\in\mathcal{S}\}$ corresponds to a stationary policy for the original MDP. Under Assumption 1, any stationary policy yields a Markov chain with equilibrium state distribution $\rho_{\pi}(s)$. Therefore, the following probability converges to a probability that is independent of $s_{1}$:
$\lim_{i\to\infty}P(X_{i}=x,S_{i}=s,S_{i+1}=s^{\prime}|s_{1})=\rho_{\pi}(s)\pi(x|s)\bm{T}(s^{\prime}|s,x)$
@@ Appendix C System Explanation and Algorithm Details
  • A quantizer then generates the codeword $\bm{U}\in\mathcal{X}^{|\mathcal{S}|\times\frac{k}{R}}$, where each element of the resultant codeword, $x_{t}=U[s_{t},t]$, represents the selected action for state $s_{t}$ at time step $t$.
  • Given the action and state, EAS channel then returns next states.
@@ C.1 Quantizer
The quantizer, $\mathcal{Q}:\mathbb{R}^{|\mathcal{S}|\times\frac{k}{R}}\rightarrow\mathcal{X}^{|\mathcal{S}|\times\frac{k}{R}}$, is designed to convert continuous coding results $|\mathcal{X}|\cdot\text{Sigmoid}(\bm{Z})$ into finite values corresponding to actions within the channel input alphabet set $\mathcal{X}=\{0,1,\ldots,|\mathcal{X}|-1\}$. For example, if there exist $5$ actions and $1$ states, then each element of the coding result $|\mathcal{X}|\cdot\text{Sigmoid}(\bm{Z})\in(0,5)$ will be rounded down into the nearest action index. If $|\mathcal{X}|\cdot\text{Sigmoid}(\bm{Z})=[1.5,0.8,2.1,3.2,4.8]$, it will be rounded down to $[1,0,2,3,4]$ as the channel input via the quantizer $\mathcal{Q}$.
@@ C.2 Critic Network and Iterative training
This process iteratively alternates between freezing one component while optimizing the other. From the experiment, we observe that the training process is more stable when the encoder is updated once for every two updates of the decoder. The effectiveness of this approach is visualized in a Fig. 11 and Table. 5. The detailed training algorithm is presented in the Algorithm. 7
@@ C.3 Training and inference algorithm
To be more specific for the training phase, we train the encoder first and then train the decoder. The critic network, with control loss, is applied to the encoder. The decoder only considers the communication loss. After training, the critic network is removed, as shown in Fig. 7 (c). Only the encoder and decoder are deployed for communication and control, as implemented in Fig. LABEL:Act_fig_implement.
@@ C.5 Vector Representation of States and Actions
This paper focuses on MDPs with finite state and action spaces, which allows us to represent states and actions as scalar integers by indexing them. This means that even when states and actions are vectors, they can be mapped to scalars. We adopt this approach in the proposed Act2Comm for ease of presentation. However, it is worth noting that using vector representations for states and actions can be beneficial in MDPs. This can be implemented with minor adjustments to the feedback structure for the encoder and the input structure for the decoder.
@@ D.1 Experiment setup
For the Act2Comm scheme, we train the model with a batch size of $4096$, a learning rate of $0.001$, and an Adam-based lookahead optimizer (Zhang et al., 2019). The inner-training step for the critic network is $s_{in}=20$, with a noise variance of $\sigma_{w}^{2}=0.1$. Each block has a length of $\mu=3$, and temperature parameter is as $\gamma=10$, $\gamma=50$, $\gamma=100$, $\gamma=200$. The performance presented is averaged over $20,000$ execution times. To investigate the trade-offs, we train the Act2Comm model with $\lambda\in[0.01,20]$.
@@ D.3 Catch the Ball
The ‘Erratic robot’ game takes place on a $4\times 4$ grid map with $16$ states and $5$ actions, as shown in Fig. 9(c). The robot is designed to minimize the number of steps required to collect goods while avoiding obstacle points. Specifically, its primary objective is to continuously take goods from designated destination points, earning a reward of $+5$ for each successful collection. The grid also includes four obstacle points, each incurring a penalty of $-2$ when encountered. The robot has five available actions: move left, move right, move up, move down, or remain stationary. Due to the instability, any movement of the robot has a probability $p\in[0,1)$ of resulting in an additional unintended step. Actions that would move the robot outside the grid boundaries always fail. In this experiment, the probability of additional operation is set to $p=0.2$. The results are presented in the Fig. 10
@@ E.1 Ablation study over different mechanisms
Additionally, to further validate the performance of our Critic network, we plot the estimated loss values generated by the Critic alongside the true loss during training. As shown in Fig. 11, the Critic network maintains accurate loss estimation throughout the training phase.
@@ E.2 Ablation study over the feedback design
Act2Comm supports both history-dependent coding, which encodes both message and feedback blocks, and Markov coding, which encodes only the message block. We examine Act2Comm scheme with both History-dependent codes and Markov codes across various coding rates $R$ for a pure communication optimization.
The experimental results are depicted in Fig. 12(a). It is evident that the BER performance improves significantly as the channel coding rate decreases. Specifically, with a coding rate $R=1/6$, the BER reaches $10^{-6}$ level, indicating a significant enhancement in performance due to the increased number of channel uses. Comparing history-dependent codes with Markov codes reveals that incorporating feedback blocks can yield substantial performance improvements. Although, from a theoretical perspective, channel feedback does not enhance the capacity of a memoryless channel, the observed benefits may be attributed to the attention mechanism applied to historical states and their corresponding actions. This mechanism effectively utilizes prior coding experience to simplify the subsequent coding process.
Figure 12: Left: Performance for different coding scheme. Right: Approaching a given policy $\bm{\bar{\pi}}$.
@@ E.4 Ablation study over a specific policy
In previous part of this paper, we focused solely on the optimal strategy as the target policy for training, aiming to optimize the control objective. However, our approach is versatile and can accommodate a broad spectrum of policies—ranging from optimal to specialized strategies—enabling users to employ the system to approach specific policies for their own tasks.
In the lucky wheel environment, now we consider a target policy $\bm{\bar{\pi}}$ given as:
@@
Notably, our method can achieve good communication performance, approximately $4\cdot 10^{-1}$, $4\cdot 10^{-2}$, and $3\cdot 10^{-3}$ for $R=1/3$, $R=1/4$, and $R=1/5$ respectively, with almost no loss in control performance. This is due to the stochastic nature of our target policy, as opposed to a deterministic one, allowing our method to adaptively learn a similar policy that maintains comparable rewards while being more favorable for channel coding. In summary, the experimental results demonstrate that our Act2Comm framework achieves satisfactory communication performance while maintaining the reward, as defined by a specific policy, at a certain level in scenarios such as those involving stochastic policies.

### v2

@@ 6 Experimental Results
We evaluate Act2Comm across three distinct MDP environments, as detailed in Appendix D, with communication performance measured by the bit error rate (BER). Due to page limitations, the results of the third environment, “Erratic robot”, are provided in the Appendix D.4.
Experiment 1: Lucky Wheel. In this game, the agent keeps spinning a wheel to accumulate rewards by choosing either clockwise or counterclockwise direction. It is modeled as an MDP with 3 states and 2 actions, the details of the environment and experimental setting are provided in Appendix D.
We examine the trade-off among the three performance metrics in Fig. 3. We set the optimal reward-maximizing policy as the target policy, and consider different code rates. By adjusting $\lambda$, we can control how closely the coding policy approximates the target policy. The shaded regions in figures (a)-(c) represent achievable regions of Act2Comm with various $\lambda$. When $\lambda$ is large, regardless of the coding rate, Act2Comm learns a policy that mirrors the target policy. In this case, all messages are mapped to the same sequence of decision rules since the target policy is stationary and deterministic. As a result, the BER is 0.5, indicating zero communication capability.
@@
Experiment 2: Catch the Ball. We next evaluate Act2Comm in “Catch the Ball”, which is an MDP with 27 states and 3 actions. This MDP includes a parameter $p$ that influences its transition matrix (see Appendix D for details). When $p=0$, the action-state channel is perfect as each action can be reliably inferred from the resulting state transition, but it becomes noisy for $p>0$.
@@ 7 Conclusion
Acknowledgments
We acknowledge funding from the UKRI for the projects AI-R (ERC Consolidator Grant, EP/X030806/1) and INFORMED-AI (EP/Y028732/1), as well as the SNS JU project 6G-GOALS under the EU’s Horizon program (Grant Agreement No. 101139232).
@@ Appendix A Notation and Definitions
To bridge the RL and FSC areas, we list and explain the notations used in this paper. Note that lowercase and uppercase bold letters represent vectors and matrices, respectively.
@@
$\mathcal{X}$, $\mathcal{Y}$, $\mathcal{S}$ Channel input, output, and state alphabets
@@ B.1 Proof of Theorem 1
    $r(s_{i},q_{i}(s_{i}))=\sum_{x_{i}\in\mathcal{X}}q_{i}(x_{i}|s_{i})\sum_{s_{i+1}\in\mathcal{S}}\bm{T}(s_{i+1}|s_{i},x_{i})\log\frac{\bm{T}(s_{i+1}|s_{i},x_{i})}{\sum_{x^{\prime}_{i}}\bm{T}(s_{i+1}|s_{i},x^{\prime}_{i}){q_{i}}(x^{\prime}_{i}|s_{i})}.$
@@
$C=\max_{\{q_{i}\}_{i\geq 1}}\lim_{N\to\infty}\frac{1}{N}\sum_{i=1}^{N}\sum_{s_{i}\in\mathcal{S}}P(s_{i})r(s_{i},q_{i}(s_{i})).$ (16)
@@
    Any stationary policy yields a Markov chain with one ergodic class and a possibly empty set of transient states (by our assumption).
@@
$C=\max_{\{\pi(\cdot|s)\in\Delta(\mathcal{X}):s\in\mathcal{S}\}}\lim_{N\to\infty}\frac{1}{N}\sum_{i=1}^{N}I(X_{i};S_{i+1}|S_{i},s_{1}).$ (17)
Note that each $\{\pi(\cdot|s)\in\Delta(\mathcal{X}):s\in\mathcal{S}\}$ corresponds to a stationary policy for the original MDP. Under the assumption that the MDP is unichain, any stationary policy yields a Markov chain with equilibrium state distribution $\rho_{\pi}(s)$. Therefore, the following probability converges to a probability that is independent of $s_{1}$:
$\lim_{i\to\infty}P(X_{i}=x,S_{i}=s,S_{i+1}=s^{\prime}|s_{1})=\rho_{\pi}(s)\pi(x|s)\bm{T}(s^{\prime}|s,x).$
@@ Appendix C System Explanation and Algorithm Details
  • A quantizer then generates the codeword $\bm{U}\in\mathcal{X}^{|\mathcal{S}|\times\frac{k}{R}}$, where each element of the resultant codeword, $x_{t}=\bm{U}[s_{t},t]$, represents the selected action for state $s_{t}$ at time step $t$.
  • Given the action and state, EAS channel then returns the next state.
@@ C.1 Quantizer
The quantizer $\mathcal{Q}:\mathbb{R}^{|\mathcal{S}|\times\frac{k}{R}}\rightarrow\mathcal{X}^{|\mathcal{S}|\times\frac{k}{R}}$ is designed to convert real-valued coding results $|\mathcal{X}|\cdot\text{Sigmoid}(\bm{Z})$ into integers corresponding to actions within the channel input alphabet set $\mathcal{X}=\{0,1,\ldots,|\mathcal{X}|-1\}$. For example, if there exist $5$ actions and $1$ state, then each element of the coding result $|\mathcal{X}|\cdot\text{Sigmoid}(\bm{Z})\in(0,5)$ will be rounded down into the nearest action index. If $|\mathcal{X}|\cdot\text{Sigmoid}(\bm{Z})=[1.5,0.8,2.1,3.2,4.8]$, it will be rounded down to $[1,0,2,3,4]$ as the channel input via the quantizer $\mathcal{Q}$.
@@ C.2 Critic Network and Iterative training
This process iteratively alternates between freezing one component while optimizing the other. From the experiment, we observe that the training process is more stable when the encoder is updated once for every two updates of the decoder. The effectiveness of this approach is visualized in a Fig. 11 and Table 5. The detailed training algorithm is presented in the Algorithm. 7
@@ C.3 Training and inference algorithm
To be more specific for the training phase, we train the encoder first and then train the decoder. The critic network, with control loss, is applied to the encoder. The decoder only considers the communication loss. After training, the critic network is removed, as shown in Fig. 7 (c). Only the encoder and decoder are deployed for communication and control, as shown in Fig. 8(b).
@@ C.5 Vector Representation of States and Actions
This paper focuses on MDPs with finite state and action spaces, which allows us to represent states and actions as scalar integers by indexing them. This means that even when states and actions are vectors, they can be mapped to scalars. We adopt this approach in the proposed Act2Comm for ease of presentation. However, it is worth noting that using vector representations for states and actions can be beneficial in some MDPs. This can be implemented with minor adjustments to the feedback structure for the encoder and the input structure for the decoder.
@@ D.1 Experiment setup
For the Act2Comm scheme, we train the model with a batch size of $4096$, a learning rate of $0.001$, and an Adam-based lookahead optimizer (Zhang et al., 2019). The inner-training for the critic network consists of $s_{in}=20$ steps, with a noise variance of $\sigma_{w}^{2}=0.1$. Each block has a length of $\mu=3$, and temperature parameter is as $\gamma=10$, $\gamma=50$, $\gamma=100$, $\gamma=200$. The performance presented is averaged over $20,000$ execution times. To investigate the trade-offs, we train the Act2Comm model with $\lambda\in[0.01,20]$.
@@ D.3 Catch the Ball
The ‘Erratic robot’ game takes place on a $4\times 4$ grid map with $16$ states and $5$ actions, as shown in Fig. 9(c). The robot is designed to minimize the number of steps required to collect goods while avoiding obstacle points. Specifically, its primary objective is to continuously take goods from designated destination points, earning a reward of $+5$ for each successful collection. The grid also includes four obstacle points, each incurring a penalty of $-2$ when encountered. The robot has five available actions: move left, move right, move up, move down, or remain stationary. Due to the instability, any movement of the robot has a probability $p\in[0,1)$ of resulting in an additional unintended step. Actions that would move the robot outside the grid boundaries always fail. In this experiment, the probability of additional operation is set to $p=0.2$. The results are presented in the Fig. 10.
@@ E.1 Ablation study over different mechanisms
Additionally, to further validate the performance of the Critic Network, we compare the estimated loss values generated by the Critic with the true loss during training. As shown in Fig. 11, the Critic Network maintains accurate loss estimation throughout the training phase.
@@ E.2 Ablation study over the feedback design
Act2Comm supports both history-dependent coding, which encodes using both message and feedback blocks, and Markov coding, which encodes using only the message block. We examine Act2Comm scheme with both history-dependent codes and Markov codes across various coding rates $R$ for a pure communication optimization.
The experimental results are depicted in Fig. 12(a). It is evident that the BER performance improves significantly as the channel coding rate decreases. Specifically, with a coding rate $R=1/6$, the BER reaches $10^{-6}$ level, indicating a significant enhancement in performance due to the increased number of channel uses. Comparing history-dependent codes with Markov codes reveals that incorporating feedback blocks can yield substantial performance improvements. Although, from a theoretical perspective, channel feedback does not enhance the capacity of a memoryless channel, it is beneficial for finite block-length coding. The observed benefits may be attributed to the attention mechanism applied to historical states and their corresponding actions. This mechanism effectively utilizes prior coding experience to simplify the subsequent coding process.
Figure 12: (a) Performance for different coding schemes. (b) Approaching a given policy $\bm{\bar{\pi}}$.
@@ E.4 Ablation study over a specific policy
In the previous experiments, we used the optimal control policy as the target policy for training the encoder, as this is better for achieving a higher control reward. However, our approach is highly flexible and can be adapted to any control policy, allowing users to tailor the system to align with specific policies for their own tasks.
In the lucky wheel environment, now we consider a sub-optimal policy $\bm{\bar{\pi}}$ as the target policy, given as:
@@
Notably, our method can achieve good communication performance, approximately $4\times 10^{-1}$, $4\times 10^{-2}$, and $3\times 10^{-3}$ for $R=1/3$, $R=1/4$, and $R=1/5$ respectively, with almost no loss in control performance. This is due to the stochastic nature of our target policy, as opposed to a deterministic one, allowing our method to adaptively learn a similar policy that maintains comparable rewards while being more favorable for channel coding. In summary, the experimental results demonstrate that our Act2Comm framework achieves satisfactory communication performance while maintaining the reward, as defined by a specific policy, at a certain level in scenarios such as those involving stochastic policies.